Three orces acting on an object are given by Fi 1.60 + 1.60 j N, F, 4.70 İ 3.30 j ) N and F3 (150 і 10j N. The object expenences an acceleration of magnitude 3 5 m/s (a) What is the direction of the acceleration? (b) What is the mass of the object? (c) If the object is initially at rest, what is its speed after 19.0 s? (d) What are the velocity components of the object after 19.0 s? 。(from the positive x axis) kg m/s CL j) m/s
Answer #1

net force acting along x axis

Fx = 1.6 - 4.7 - 45 = - 48.1 N

net force acting along y axis

Fy = 1.6 - 3.3 - 43 = - 44.7 N

net force

F^2 = Fx^2 + Fy^2

F = 65.66 N

============

a)

Direction of Acceleration

phi = 180 + arctan ( 44.7/ 48.1)

phi = 222.9 ccw from +x axis

=========

b)

F = ma

m = F/a = 65.66/ 3.55 = 18.5 kg

========

c)

Vx = ( Fx/ m) t = - 48.1/ 18.5 * 19 = - 49.41 m/s

Vy = ( Fy/m) t = - 44.7/18.5 *19 = - 45.91 m/s

net velocity

V^2 = Vx^2 + Vy^2

V = 67.45 m/s

=========

d)

V = - 49.41 i + ( - 45.91) j

========

do comment in case any doubt, will reply for sure . Goodluck